OpenStudy (anonymous):

3 min. 5 sec. convert into hours in terms of fraction

OpenStudy (anonymous):

(3/60)+(5/3600)=185/3600.... yu can still reduce it

OpenStudy (anonymous):

3 min 5 sec = (3*60 + 5) sec = 185 sec 3600 sec = 1 hr 185 ------- h 3600 37 --- h 720